OpenStudy (abbycross167):

30. Which numbers are solutions of the equation | x – 7 | + 5 = 17? a. -19 and 15 b. -15 and 19 c. -5 and 19 d. -15 and 29 My Answer~ c

OpenStudy (anonymous):

easy enough check right?\[|-5-7|+5=|-12|+5=12+5=17\] so that one is right

OpenStudy (anonymous):

and 19 works as well as usual, it is C

OpenStudy (abbycross167):

Okay, thank you so much!!

OpenStudy (anonymous):

lol you did all the work, but you are welcome just the same

OpenStudy (danjs):

if they gave you choices, you can know nothing and still test each answer and get it right

OpenStudy (abbycross167):

Oh ok, thank you as well @DanJS I really appreciate y'all's help